Thirty-five percent of all adults read at least one newspaper daily. Suppose you select a sample of 501 adults. The sample proportion is the proportion of adults in this sample who read at least one newspaper daily.

The mean of the sampling distribution of this sample proportion is: Enter you answer; mean

The standard deviation of the sampling distribution of this sample proportion, rounded to four decimal places, is:
Answer #1

Solution

Given that,

p = 0.35

1 - p = 1 - 0.35 = 0.65

n = 501

\mu\hat p = p = 0.35

\sigma\hat p =  \sqrt[p ( 1 - p ) / n] =  \sqrt [(0.35 * 0.65) / 501 ] = 0.0213